Calculate the solubility of NiCO3 in moles per liter and grams per liter (Ksp = 1.4 x 10-7)


1
Expert's answer
2021-10-25T03:48:48-0400

NiCO3 --> Ni2+ + CO32-

Since 1 formula unit of NiCO3 produces 2 ions, the formula for molar solubility (s) is

"K_{sp}=s^2"

Therefore,

"s=\sqrt{K_{sp}}=\sqrt{1.4\times10^{-7}}=3.74\times10^{-4}\space{mol/L}"

To find the solubility in grams per liter, s should be multiplied by the molar mass:

"3.74\times10^{-4}\frac{mol}{L}\times118.7\frac{g}{mol}=0.0444\space{g/L}"


Answer: "3.74\times10^{-4}\space{mol/L}" ; 0.0444 g/L